# Initialization of Monocular Visual Navigation for Autonomous Agents Using Modified Structure from Small Motion

### v1

@@ II-C Step 2: Translation and Inverse Depth Estimation
We proceed by fixing the rotations $\bm{\mathrm{\theta}}_{i}^{*}$ as estimated in section II-B and by initializing the translations $\bm{\mathrm{r}}_{i}^{(0)}\leftarrow\bar{\bm{\mathrm{r}}}_{i}^{*}/\bar{w}$ and the inverse depths $w_{j}^{(0)}\leftarrow\bar{w}$. These initial estimates are then refined by a BA procedure restricted to estimating only translations and inverse depths. We leverage the Georgia Tech Smoothing And Mapping Library (GTSAM) [24] to implement and solve the restricted BA using factor graph optimization and a Levenberg Marquardt (LM) solver. A custom factor is defined to enable simultaneous optimization of both $\bm{\mathrm{r}}_{i}$ and $w_{j}$, encoding the image disparity residual.
To ensure that all landmark points remain in front of the camera (the cheirality condition), we introduce a re-parameterization of $w_{j}$ using the the soft-plus function $\mathtt{sp}(x)$ given by $\mathtt{sp}(x)={\ln\left(1+\exp\left(\alpha x\right)\right)}/{\alpha}$. To avoid numerical instability, we adopt the approach outlined in [25], such that
@@
where $\mathtt{log1p}(x)$ is an implementation of the $\ln(1+x)$ function which is numerically accurate for values of $x\rightarrow 0^{+}$.
We incorporate the soft-plus function in Eq. (8) into a custom factor $\bm{\mathrm{\varepsilon}}_{ij}^{\mathrm{step2}}$ that encodes as residual the 2D disparity between the measured image point $\bm{\mathrm{p}}_{ij}^{\mathrm{m}}$ and the expected projection of the landmark as a function of a translation $\bm{\mathrm{r}}\in\mathbb{R}^{3}$ and a landmark parameter $\omega\in\mathbb{R}$, while maintaining the rotation $\bm{\mathrm{\theta}}_{i}^{*}$ fixed, as given by
$\bm{\mathrm{\varepsilon}}_{ij}^{\mathrm{step2}}(\bm{\mathrm{r}},$ $\,\omega)=$
$\bm{\mathrm{p}}_{ij}^{\mathrm{m}}-\left\langle\bm{\mathrm{K}}\left(\left(I_{3}+[\bm{\mathrm{\theta}}_{i}^{*}]_{\times}\right)\bm{\mathrm{x}}_{0j}^{\mathrm{m}}+\mathtt{sp}(\omega)\bm{\mathrm{r}}\right)\right\rangle.$ (9)
We make use of Eq. (9) to construct the restricted BA problem for poses $i=1,\ldots,n$, and landmarks $j=1,\ldots,m$, given as
@@
where $\Omega(x)$ is the robust Huber cost, $\|\bm{\mathrm{\varepsilon}}\|_{\Sigma}=\sqrt{\bm{\mathrm{\varepsilon}}^{\top}\Sigma^{-1}\bm{\mathrm{\varepsilon}}}$, and $\Sigma_{ij}$ corresponds to the covariance of the measurement $j$ taken at pose $i$.
We solve Eq. (10) using the LM iterative procedure. An initial value for the parameter $\omega_{j}^{(0)}\leftarrow\mathtt{sp}^{-1}(w_{j}^{(0)})$ is provided to the LM solver from the inverse of the soft-plus function $\mathtt{sp}^{-1}(y)$, given as $\mathtt{sp}^{-1}(y)={\ln\left(\exp\left(\alpha y\right)-1\right)}/{\alpha}$.
The optimized $\omega_{j}^{*}$ values are accordingly converted to $w_{j}^{*}$ using Eq. (8), guaranteeing positive inverse depths. The minimizers ($\left(\{\bm{\mathrm{r}}_{i}^{*}\}_{i=1}^{n},\{w_{j}^{*}\}_{j=1}^{m}\right)$ resulting from Eq. (10) may now be used to warm-start the initial values of step 3. The combination of our step 1 and step 2 procedures avoid the solution ambiguity that arises in pose estimation using essential matrix decomposition [14].
@@ II-D Step 3: Full Bundle Adjustment
Equipped with estimates for all parameters $\bm{\mathrm{\theta}}_{i}^{*}$, $\bm{\mathrm{r}}_{i}^{*}$, and $w_{j}^{*}$, we now solve a full BA to refine these variables and compute the camera trajectory and the coordinates of the landmarks.
In contrast with Sections II-B and II-C, the step 3 parameterization of the expected landmark coordinates does not involve the measurement of the landmark in the reference camera frame.
@@
Instead, we introduce variables to allow for in-camera-plane variation of the landmark coordinate vector to account for quantization of the image feature coordinate measurement. Specifically, we adopt the inverse depth approach in [26], Fig. 1, which re-parameterizes the 3D point $\bm{\mathrm{y}}_{0j}$ such that
$\bm{\mathrm{y}}_{0j}=\bm{\mathrm{m}}_{j}/\rho_{j},$
where $\bm{\mathrm{m}}_{j}=\bm{\mathrm{m}}(\psi_{j},\phi_{j})$ is a unit directional vector from the reference camera frame to the $j$-th landmark, with $\bm{\mathrm{m}}(\psi,\phi)$ given by
$\bm{\mathrm{m}}(\psi,\phi)=[\cos{\phi}\sin{\psi},-\sin{\phi},\cos{\phi}\cos{\psi}]^{\top},$
and where $\rho_{j}$ is the inverse depth, given as $\rho_{j}=1/\|\bm{\mathrm{y}}_{0j}\|$. [^1] We note that $\rho_{j}=1/\|\bm{\mathrm{y}}_{0j}\|$ replaces the previous definition of the inverse depth $w_{j}=1/Z_{j}$. Accordingly, $\bm{\mathrm{m}}_{j}$ is parameterized by the azimuth angle $\psi_{j}$ and the elevation angle $\phi_{j}$, which, for $\bm{\mathrm{y}}_{0j}=[X_{j},Y_{j},Z_{j}]^{\top}$, can be computed using $\psi_{j}=\arctan(X_{j},Z_{j})$, and $\phi_{j}=\arctan(-Y_{j},\sqrt{X_{j}^{2}+Z_{j}^{2}})$.
As in Section II-C, we make use of the soft-plus function re-parameterization to guarantee the strict positiveness of $\rho_{j}$, and thus $\rho_{j}=\mathtt{sp}(\omega_{j})$.
We define a new custom factor $\bm{\mathrm{\varepsilon}}_{ij}^{\mathrm{step3}}$ that encodes as residual the 2D disparity between the measured image point $\bm{\mathrm{p}}_{ij}^{\mathrm{m}}$ and the expected projection of the corresponding landmark as a function of the viewing camera’s rotation $\bm{\mathrm{R}}\in\mathrm{SO}(3)$, its translation vector $\bm{\mathrm{r}}\in\mathbb{R}^{3}$, and the landmark’s parameters $\omega,\psi,\phi\in\mathbb{R}$, such that
@@
Thus, the additional DoF ($\psi_{j}$ and $\phi_{j}$) and the inverse depth $\rho_{j}$, determine the 3D location of landmark $j$. At the same time, we maintain the reference frame at a fixed pose such that $\bm{\mathrm{R}}_{0}=I_{3}$ and $\bm{\mathrm{r}}_{0}=\bm{\mathrm{0}}$. Thus, reformulating Eq. (11), for each measurement $\bm{\mathrm{x}}_{0j},\,j=1,\ldots,m$, captured in the reference frame, we constrain the variables $\psi_{j}$ and $\phi_{j}$, yielding the additional factor $\bm{\mathrm{\varepsilon}}_{0j}^{\mathrm{prior}}$, which encodes the 2D disparity between the measured image point $\bm{\mathrm{p}}_{0j}^{\mathrm{m}}$ and the expected projection of the corresponding landmark as a function of the landmark’s parameters $\psi$ and $\phi$, given by
$\bm{\mathrm{\varepsilon}}_{0j}^{\mathrm{prior}}(\psi,\phi)=\bm{\mathrm{p}}_{0j}^{\mathrm{m}}-\left\langle\bm{\mathrm{K}}\bm{\mathrm{m}}(\psi,\phi)\right\rangle.$ (12)
We make use of Eq. (11) and Eq. (12) to construct the full BA problem for poses $i=1,\ldots,n$, and landmarks $j=1,\ldots,m$, given by
@@
where $\Omega(x)$, $\|*\|_{\Sigma}$ and $\Sigma_{ij}$ are similar to their description in Section II-C.
The minimizers ($\left(\{(\bm{\mathrm{R}}_{i}^{**},\bm{\mathrm{r}}_{i}^{**})\}_{i=1}^{n},\{(\omega_{j}^{**},\psi_{j}^{**},\phi_{j}^{**})\}_{j=1}^{m}\right)$ to problem Eq. (13) are then used to initialize and constrain the ensuing full SLAM problem.
@@ III Experiments
We evaluate our initialization pipeline on 100 synthetic small-motion image sequences generated in a custom Unreal Engine 5 [27] simulation platform.
@@
Each sequence represents a weak-perspective, center-pointing inspection trajectory and consists of 20 images captured over a 2-second duration at 10 frames per second using a simulated camera with a 14.9-degree field of view (FOV). The target RSO in the image sequences is a highly detailed model of the Hubble Space Telescope (HST), positioned 100 meters from the camera.
The simulated images incorporate dynamic and harsh lighting conditions to closely mimic the challenges of visual estimation in space (see Fig. 2). The motion of the scene emulates a tumbling RSO being tracked by a chaser.
Figure 3: Normalized translation, rotation, and depth error results between the proposed method (blue) and USAC_FM_8PTS on trajectories spanning from 0 to 6 degrees of parallax. The USAC_FM_8PTS approach failed to reach a success state in any of the attempted sequences.
We compare the proposed method to the 3-step SfSM pipeline delineated in [12] and the state-of-the-art USAC_FM_8PTS algorithm [16], which is a wide-baseline, model selection approach.
@@
For consistency, we set the reference frame of the estimated trajectory to align with that of the ground truth trajectory, and we normalize the estimated translations by the magnitude of the $n$-th frame translation. Depths are scaled by the same normalizing factor to maintain a consistent scale. For each sequence, we compute (a) the absolute translation error $\bm{\mathrm{\epsilon}}_{i}^{\mathrm{ATE}}=\bm{\mathrm{R}}_{i}^{\top}\bm{\mathrm{r}}_{i}-(\bm{\mathrm{R}}_{i}^{\mathrm{true}})^{\top}\bm{\mathrm{r}}_{i}^{\mathrm{true}}$ across all frames; (b) the absolute rotation error $\bm{\mathrm{\epsilon}}_{i}^{\mathrm{ARE}}=\bm{\mathrm{\theta}}_{i}-\bm{\mathrm{\theta}}_{i}^{\mathrm{true}}$ across all frames; and (c) the depth error $\bm{\mathrm{\epsilon}}_{j}^{\mathrm{depth}}=Z_{j}-Z_{j}^{\mathrm{true}}$ at the reference frame across all landmarks. Depth error is obtained from a comparison with the scale normalized depth map of the simulated RSO, shown in Fig. 4.
We compute the root-mean-square (RMS) for each error evalue, as given by
@@
where $N$ is the sample size and $*\in\{\mathrm{ATE},\,\mathrm{ARE},\,\mathrm{depth}\}$. To represent data trends across all sequences, we compute the mean of the aforementioned RMS errors.
@@
Proposed 57.6 0.222 0.027 0.587 0.039 - 0.149 - 49.2
Ha et al. [12] 38.8 0.442 0.136 1.08 0.106 - 0.185 - 19.6
USAC FM [16] 0.0 0.970 0.046 4.11 0.035
All evaluated methods were implemented in C++ using the GTSAM library and OpenCV [17]. We provided the pipelines with the same inlier ORB [28] feature tracks from a shared front-end. Furthermore, both SfSM methods used the same LM optimizer parameters. The experiments were conducted on a machine equipped with an AMD Ryzen 7800X3D CPU and 32 GB of DDR5 RAM.
The results, summarized in Table I, reveal that the proposed method is significantly more robust than the method in [12] or the state-of-the-art USAC_FM_8PTS method when initializing from weak-perspective, center-pointing motions with dominant planes in the scene (HST’s solar panels).
The proposed method is able to successfully initialize in 57.6% of the sequences, an improvement of 48.5% over the original 3-step approach, which could only initialize in 38.8% of the sequences. The USAC_FM_8PT method failed to successfully initialize in any of the tested sequences.
In Fig. 5, the smaller height of the box plots for our method indicates a lower spread in error across all SLAM variables, demonstrating more reliable and robust performance. Of note is the narrower interquartile range in rotation, which demonstrates improved small motion rotation estimation compared to [12].
Figure 5: Error value distribution, median (blue line), and mean (red circle) for scale normalized translations (left), rotations (middle), and scale normalized depths (right) for the proposed method and the method in [12].
Moreover, the proposed method is, on average, 49.7% more accurate in translation estimation, 80.1% more accurate in rotation estimation, and 45.6% more accurate in depth estimation. Thus, we are able to more effectively and consistently overcome the bas-relief ambiguity while attempting to initialize in small-motion image sequences, as seen in Fig. 6.
@@
Figure 6: The SLAM solution one step after initialization, where the estimated poses (RGB, larger triads) are close to the ground truth poses (CYM, smaller triads) and the estimated landmarks (cyan) are coincident with the ground truth HST object.
Although more robust and accurate, our algorithm demonstrates improvements obtained at the cost of an increased step 3 computation time of 2.5 times that of the competing method. The inverse depth and 3D point re-parameterizations in step 3 enhance the robustness of the solution, but also introduce additional variables to the the optimization, thus increasing the dimensionality of the search space. A key challenge arises from the vanishing gradient issue, which hinders convergence [29]. The problem is exacerbated by ambiguous observations resulting from small apparent motion and scale differences between the optimization variables. Yet, when considering the scale and range of SLAM trajectories in space, the added computation time becomes less significant.

### v2

@@ II-C Step 2: Translation and Inverse Depth Estimation
We proceed by fixing the rotations $\bm{\mathrm{\theta}}_{i}^{*}$, estimated in section II-B, and initializing the translations as $\bm{\mathrm{r}}_{i}^{(0)}\leftarrow\bar{\bm{\mathrm{r}}}_{i}^{*}/\bar{w}$ and the inverse depths as $w_{j}^{(0)}\leftarrow\bar{w}$. These initial estimates are then refined by a BA procedure restricted to estimating only translations and inverse depths. We utilize the Georgia Tech Smoothing And Mapping Library (GTSAM) [23] to implement and solve the restricted BA via factor graph optimization and a Levenberg Marquardt (LM) solver. A custom factor is defined to enable simultaneous optimization of both $\bm{\mathrm{r}}_{i}$ and $w_{j}$, encoding the image disparity residual.
To ensure all landmarks satisfy the cheirality condition, we reparameterize $w_{j}$ using the soft-plus function $\mathtt{sp}(x)={\ln\left(1+\exp\left(\alpha x\right)\right)}/{\alpha}$, such that $w_{j}=\mathtt{sp}(\omega_{j})$, where $\omega_{j}\in\mathbb{R}$ is now the landmark-related optimization variable, initialized with $\omega^{(0)}_{j}\leftarrow\mathtt{sp}^{-1}(w^{(0)}_{j})$. To avoid numerical instability, by use the equation from [24], we have that
@@
where $\mathtt{log1p}(x)$ is a numerically stable implementation of $\ln(1+x)$, particularly for $x\rightarrow 0^{+}$.
We incorporate Eq. (8) into a custom factor $\bm{\mathrm{\varepsilon}}_{ij}^{\mathrm{step2}}$ that encodes, as residual, the 2D disparity between the measured image point $\bm{\mathrm{p}}_{ij}^{\mathrm{m}}$ and the expected projection of the corresponding landmark. The residual is modeled as a function of the parameters $\bm{\mathrm{r}}_{i}$ and $\omega_{j}$, while keeping $\bm{\mathrm{\theta}}_{i}^{*}$ fixed, as in
$\bm{\mathrm{\varepsilon}}_{ij}^{\mathrm{step2}}(\bm{\mathrm{r_{i}}},$ $\,\omega_{j})=$
$\bm{\mathrm{p}}_{ij}^{\mathrm{m}}-\left\langle\bm{\mathrm{K}}\left(\left(I_{3}+[\bm{\mathrm{\theta}}_{i}^{*}]_{\times}\right)\bm{\mathrm{x}}_{0j}^{\mathrm{m}}+\mathtt{sp}(\omega_{j})\bm{\mathrm{r_{i}}}\right)\right\rangle.$ (9)
We apply Eq. (9) to construct the restricted BA problem for poses $i=1,\ldots,n$ and landmarks $j=1,\ldots,m$, as in
@@
where $\Omega(x)$ is the robust Huber cost, $\|\bm{\mathrm{\varepsilon}}\|_{\Sigma}=\sqrt{\bm{\mathrm{\varepsilon}}^{\top}\Sigma^{-1}\bm{\mathrm{\varepsilon}}}$, and $\Sigma_{ij}$ is the covariance of the measurement $j$ taken at pose $i$. We solve Eq. (10) using the LM iterative procedure.
The optimized $\omega_{j}^{*}$ values are converted to $w_{j}^{*}$ using Eq. (8), ensuring positive inverse depths. The minimizers ($\left(\{\bm{\mathrm{r}}_{i}^{*}\}_{i=1}^{n},\{w_{j}^{*}\}_{j=1}^{m}\right)$ resulting from Eq. (10) are used to initialize the procedure in Section II-D. The procedures from Sections II-B and II-C collectively avoid the solution ambiguity associated with pose estimation from essential matrix decomposition [13].
@@ II-D Step 3: Full Bundle Adjustment
Equipped with estimates for all parameters $\bm{\mathrm{\theta}}_{i}^{*}$, $\bm{\mathrm{r}}_{i}^{*}$, and $w_{j}^{*}$, we solve a full BA to compute the camera trajectory and the coordinates of the landmarks.
@@
Deviating from the landmark definition in Sections II-B and II-C, we reformulate the parameterization of the expected landmark coordinates to account for image feature quantization by introducing variables for in-camera-plane variation of the coordinates instead of directly exploiting the reference camera frame’s image feature measurement. We adopt the inverse depth approach in [25], Fig. 1, re-parameterizing the 3D point $\bm{\mathrm{y}}_{0j}$ as $\bm{\mathrm{y}}_{0j}=\bm{\mathrm{m}}_{j}/\rho_{j}$, where $\bm{\mathrm{m}}_{j}=\bm{\mathrm{m}}(\psi_{j},\phi_{j})$ is a unit directional vector from the reference camera frame to the $j$-th landmark. Here, $\bm{\mathrm{m}}(\psi,\phi)$ is defined as $\bm{\mathrm{m}}(\psi,\phi)=[\cos{\phi}\sin{\psi},-\sin{\phi},\cos{\phi}\cos{\psi}]^{\top}$, and $\rho_{j}=1/\|\bm{\mathrm{y}}_{0j}\|$ is the inverse depth, replacing the prior definition $w_{j}=1/Z_{j}$. Accordingly, $\bm{\mathrm{m}}_{j}$ is parameterized by the azimuth angle $\psi_{j}$ and the elevation angle $\phi_{j}$, which, for $\bm{\mathrm{y}}_{0j}=[X_{j},Y_{j},Z_{j}]^{\top}$, are computed as $\psi_{j}=\arctan(X_{j}/Z_{j})$ and $\phi_{j}=\arctan(-Y_{j}/\sqrt{X_{j}^{2}+Z_{j}^{2}})$. As in Section II-C, we apply Eq. (8) in $\rho_{j}=\mathtt{sp}(\omega_{j})$ to ensure $\rho_{j}$ remains strictly positive, guaranteeing the cheirality condition.
We define a custom factor $\bm{\mathrm{\varepsilon}}_{ij}^{\mathrm{step3}}$ that encodes the residual described in Section II-C; however, the residual is now modeled as a function of the camera’s rotation $\bm{\mathrm{R}}\in\mathrm{SO}(3)$, the camera’s translation $\bm{\mathrm{r}}\in\mathbb{R}^{3}$, and the landmark parameters $\omega,\psi,\phi\in\mathbb{R}$, such that
@@
Thus, $\rho_{j}$, $\psi_{j}$, and $\phi_{j}$ determine the 3D location of landmark $\bm{\mathrm{y}}_{0j}$.
Meanwhile, we maintain the reference frame at a fixed pose with $\bm{\mathrm{R}}_{0}=I_{3}$ and $\bm{\mathrm{r}}_{0}=\bm{\mathrm{0}}$. Reformulating Eq. (11), for each reference frame measurement $\bm{\mathrm{x}}_{0j},\,j=1,\ldots,m$, we constrain $\psi_{j}$ and $\phi_{j}$ to yield an additional factor $\bm{\mathrm{\varepsilon}}_{0j}^{\mathrm{prior}}$. The factor encodes the residual described in Section II-C, but now as a function of $\psi_{j}$ and $\phi_{j}$, as in
$\bm{\mathrm{\varepsilon}}_{0j}^{\mathrm{prior}}(\psi_{j},\phi_{j})=\bm{\mathrm{p}}_{0j}^{\mathrm{m}}-\left\langle\bm{\mathrm{K}}\bm{\mathrm{m}}(\psi_{j},\phi_{j})\right\rangle.$ (12)
We use Eq. (11) and Eq. (12) to construct the BA problem for poses $i=1,\ldots,n$ and landmarks $j=1,\ldots,m$, as in
@@
where $\Omega(x)$, $\|*\|_{\Sigma}$ and $\Sigma_{ij}$ are described in Section II-C. The minimizers ($\left(\{(\bm{\mathrm{R}}_{i}^{**},\bm{\mathrm{r}}_{i}^{**})\}_{i=1}^{n},\{(\omega_{j}^{**},\psi_{j}^{**},\phi_{j}^{**})\}_{j=1}^{m}\right)$ to problem Eq. (13) serve as the vSLAM initialization solution.
@@ III Experiments
We evaluate our initialization pipeline on 100 synthetic small-motion image sequences generated in a custom simulation platform in Unreal Engine 5 [26].
@@
Each sequence represents a weak-perspective, center-pointing RSO inspection trajectory and consists of 20 images captured at 10 frames per second using a simulated camera with a 14.9-degree field of view. The RSO is a tumbling, highly-detailed model of the Hubble Space Telescope (HST), positioned 100 meters from the camera. The simulated images of the tumbling HST, as shown in Fig. 2, exhibit dynamic illumination conditions that closely mimic the challenges of inspection in the vicinity of a non-cooperative RSO.
@@
We compare the proposed method to the SfSM pipeline developed by Ha et al. [11] and the USAC_FM_8PTS algorithm [15]. For consistency, we align the reference frame of the estimated trajectory with that of the ground truth trajectory and we normalize the estimated translations by the magnitude of the $n$-th frame translation. Depths are scaled by the same normalizing factor to ensure a consistent scale. For each sequence, we compute (a) the absolute translation error $\bm{\mathrm{\epsilon}}_{i}^{\mathrm{ATE}}=\bm{\mathrm{R}}_{i}^{\top}\bm{\mathrm{r}}_{i}-(\bm{\mathrm{R}}_{i}^{\mathrm{true}})^{\top}\bm{\mathrm{r}}_{i}^{\mathrm{true}}$ across all frames; (b) the absolute rotation error $\bm{\mathrm{\epsilon}}_{i}^{\mathrm{ARE}}=\bm{\mathrm{\theta}}_{i}-\bm{\mathrm{\theta}}_{i}^{\mathrm{true}}$ across all frames; and (c) the depth error $\bm{\mathrm{\epsilon}}_{j}^{\mathrm{depth}}=Z_{j}-Z_{j}^{\mathrm{true}}$ at the reference frame across all landmarks. Depth error is obtained by comparing estimated depths against the scale normalized depth map of the simulated RSO, shown in Fig. 3. We compute the root-mean-square (RMS) for each error value, given by
@@
where $N$ is the sample size and $*\in\{\mathrm{ATE},\,\mathrm{ARE},\,\mathrm{depth}\}$. To represent data trends across all sequences, we compute the mean of the RMS errors.
@@
Proposed 62.0 0.277 0.032 0.605 0.040 - 0.156 - 55.8
Ha et al. [11] 39.0 0.818 0.112 1.25 0.107 - 0.192 - 21.2
USAC [15] 0.0 0.970 0.046 4.11 0.035
All evaluated methods were implemented in C++ using the GTSAM library and OpenCV [16]. We provided the pipelines with the same inlier ORB [27] feature tracks from a shared front-end. Furthermore, both SfSM methods used the identical LM optimizer parameters. The experiments were conducted on a machine equipped with an AMD Ryzen 7800X3D CPU and 32 GB of DDR5 RAM.
The results, summarized in Table I, demonstrate that the proposed method is significantly more robust than both Ha’s SfSM method and the state-of-the-art USAC_FM_8PTS algorithm when initializing from realistic inspection trajectories that emulate a tumbling RSO.
Figure 4: Error values for scale-normalized translations, rotations, and scale-normalized depths for the proposed method (blue) and the USAC_FM_8PTS algorithm [15] (red) on trajectories spanning from 0 to 6 degrees of parallax.
Figure 5: Error value distribution, median (blue line), and mean (red circle) for scale-normalized translations (left), rotations (middle), and scale-normalized depths (right) for the proposed method and the method in [11].
@@
Figure 6: SLAM solution one step after initialization, with estimated poses (RGB) closely aligned with ground truth poses (CYM) and estimated landmarks (cyan) coinciding with the ground truth HST object.
The proposed method successfully initialized in 62% of the sequences—a $59$% improvement over Ha’s approach, which only initialized in 39% of the sequences. In contrast, the USAC_FM_8PT method, tested at parallaxes up to 6 degrees, failed to initialize in any of the sequences and demonstrated lower accuracy than the proposed method, as shown in Fig. 4. In Fig. 5, the smaller height of the box plots for our method indicates a lower spread in error across all SLAM variables, demonstrating more reliable and robust performance. Moreover, compared to Ha’s approach, the proposed method is, on average, 66.1% more accurate in translation estimation, 71.4% more accurate in rotation estimation, and 51.6% more accurate in depth estimation. Thus, we are able to more effectively and consistently overcome the bas-relief ambiguity when initializing in RSO inspection trajectories, as seen in Fig. 6.
Although more robust and accurate, the proposed algorithm achieves these improvements at the cost of an increased step 3 computation time that is 1.63 times longer than the competing SfSM method. The inverse depth and landmark re-parameterizations in step 3 enhance the robustness of the solution, but also introduce additional variables to the the optimization, increasing the dimensionality of the search space. A key challenge arises from the vanishing gradient issue, which hinders convergence [28]. The problem is exacerbated by ambiguous observations and scale differences between the optimization variables. However, when considering the scale and range of RSO inspection trajectories, the added computation time is less impactful.